How do you simplify 8 and 9 sixths?
Answer by jamesdressler.com(1)

Change the 9 sixths to a mixed number and add the whole number to 8.
8 9/6 =
8 + 9/6 =
8 + (6 + 3)/6 =
8 + 6/6 + 3/6 =
8 + 1 + 3/6 =
9 + 3/6 =
9 + 1/2 =
Nine and one half.